CHAPTER 8 HW - MICRO



Use the table below to answer the next 3 questions

Units of Output

Total Fixed Cost

Total Variable Cost

1

$1000

$200

2

450

3

800

4

1350

5

1950

5. Given the cost schedule above, the ATC of producing 4 units is

(a)        $550                                                    (c)        $1000

(b)        $588                                                    (d)        $2350

7. Given the cost schedule above, the TC of producing 3 units of output is

(a)        $800                                                    (c)        $1800

(b)        $1000                                                  (d)        $2800

8. Given the cost schedule above, it can be seen that the MC of the 3rd unit produced is

(a)        $350                                                    (c)        $600

(b)        $550                                                    (d)        $800

9.   AFC is

(a)        constant at all levels of output               (c)        the difference between AVC and ATC

(b)        less than MC when MC is falling           (d)        only important in the long run

10. A firms' short run AVC cost curve is at a minimum at the quantity where

(a)        the AVC is equal to the ATC                 (c)        the AVC is equal to the AFC

(b)        the AVC is equal to the MC                  (d)        the MC is at a minimum
Answer #1

(5)

option (b) 588   

ATC = TC/number of outputs produced

= 5350/4 (TC= TFC+TVC)

= 587.5 588

7)

option (c) 1800

TC= TFC+TVC

= 1000+800

= 1800

9)

option (c) the difference between AVC and ATC